# ColaVLA: Leveraging Cognitive Latent Reasoning for Hierarchical Parallel Trajectory Planning in Autonomous Driving

### v2

@@ 1 Introduction
Autonomous driving aims to predict safe and comfortable motion from rich multimodal observations. Early systems were organised as modular perception–prediction–planning stacks with dedicated 3D perception [qi2017pointnet, lang2019pointpillars, sun2020scalability, li2022bevformer, zheng2024denseg, zheng2025densegrounding, peng2025proxytransformation], forecasting [chai2019multipath, gu2023vip3d, liu2021multimodal, gu2021densetnt], and planning components [song2020pip, ettinger2021large, caesar2021nuplan, chitta2022transfuser, liao2024diffusiondrive]. Recent end-to-end approaches unify the stack and learn from pixels to waypoints or controls in a single pipeline [hu2022stp3, jiang2023vad, hu2023planning, chen2024vadv2, zheng2024genad, weng2024drive, sun2024sparsedrive]. In parallel, vision–language models (VLMs) [gpt3, DeepSeek-VL2, bai2025qwen2, wang2023cogvlm, liu2024visual] are increasingly integrated to inject cross-modal priors and world knowledge, either as agentic planning models that output controls or trajectories [xu2024drivegpt4, mao2023language, sima2023drivelm, hwang2024emma, xing2024openemma, wang2024omnidrive] or as reasoning assistants that guide end-to-end modules [jiang2024senna, tian2024drivevlm, Chen_2025_CVPR].
Despite rapid progress, important gaps remain between accuracy in benchmarks and reliability in deployment. Modular systems provide interpretable components and strong geometric priors, but brittle interfaces can propagate errors and make global optimization difficult. End-to-end systems reduce manual interfaces and achieve high open-loop accuracy [hu2022stp3, jiang2023vad, hu2023planning, chen2024vadv2, zheng2024genad, weng2024drive], yet they often rely on sparse trajectory supervision, intertwine perception and control in ways that obscure causal structure, and struggle to generalize to out-of-distribution scenarios. Text-based VLM planners add powerful priors but introduce some practical issues: (1) Modality mismatch. Discrete text tokens do not align with the continuous geometry and dynamics of trajectories and can lead to format violations or physically inconsistent waypoints [mao2023language, hwang2024emma]. (2) Chain-of-thought reasoning latency. The computation overhead of autoregressive decoding arises from its iterative token-by-token generation, where each new token depends on previously ones, causing the sequence to grow over time and significantly increasing inference delay.
@@ 2.1 End-to-end Autonomous Driving
Classical autonomous driving stacks decompose perception, prediction, and planning into separate modules, easing engineering but causing information loss and suboptimal coordination. End-to-end (E2E) systems instead learn a single differentiable mapping from sensor inputs to trajectories or control commands. Early demonstrations proved feasibility but suffered from limited transparency and data efficiency. Recent works enhance this paradigm with richer supervision and unified architectures: UniAD [hu2023planning] integrates perception and planning within a Transformer, and VAD [jiang2023vad, chen2024vadv2] employs vectorized abstractions for efficiency and uncertainty modeling. Other directions explore task dependency [weng2024drive], latent generative modeling [zheng2024genad, liao2024diffusiondrive], or simplified ego-centric controllers [admlp, li2024ego]. Despite progress, most E2E methods rely on sparse supervision from a single ground truth trajectory, which limits reasoning and diversity, highlighting the need for higher-level semantic and cognitive guidance.
@@ 2.2 VLMs for Autonomous Driving
Vision–Language Models (VLMs) bring broad world knowledge and remarkable reasoning ability to autonomous driving, inspiring two main integration paradigms. Single-system approaches treat planning as language modeling: visual inputs and prompts are translated into textual trajectories or control commands with chain-of-thought explanations [hwang2024emma, mao2023gpt, xing2024openemma, xu2024drivegpt4, chen2024driving]. Dual-system designs pair a VLM with an end-to-end planner, where the VLM proposes high-level commands or low-frequency trajectories to guide a lightweight action controller [tian2024drivevlm, jiang2024senna, Chen_2025_CVPR]. Other efforts align vision–language features in BEV [wang2024drive], extend reasoning to 3D [wang2024omnidrive], or use VLMs as teachers to supervise E2E planning models [xu2024vlm]. While these methods improve interpretability and generalization, they still rely on explicit text-level reasoning or suffer from a feature gap between the VLM and the downstream action planner. This motivates our unified vision–language–action framework, which performs reasoning entirely in a shared latent space, bridging perception, cognition, and control within a single coherent architecture.
@@ 3.1 Framework Overview
Given multimodal representation $\mathbf{S}_{t}$ comprising multi–view images, LiDAR points, radar grids, a text prompt and ego state, or a subset thereof, the trajectory planning task reduces to mapping $\mathbf{S}_{t}$ to a $K$-step trajectory $\widehat{\mathbf{Y}}_{t}=[(x_{t+1},y_{t+1}),\ldots,(x_{t+K},y_{t+K})]$. Existing planning systems generally consist of two unified learnable components: Reasoner ($\mathcal{R}_{\theta}$), usually implemented as ResNet [he2016resnet], PointNet [qi2017pointnet], Q-Former [dalmonte2025qformerautoencodermodernframework], or VLM Transformer that extracts and fuses multimodal features or optionally refines them with chain-of-thought, and Planner ($\mathcal{P}_{\phi}$), usually instantiated as an deterministic MLP planner, stochastic diffusion sampler, or standard Transformer [attention] decoder that injects action queries and regresses continuous way-points. Their interaction is succinctly captured by
@@ 3.2 Cognitive Latent Reasoning
Recent vision–language models show strong generalization in autonomous driving by performing a driver-like reasoning process: perceiving the scene, identifying critical entities, rethinking carefully and deciding final strategies. However, most existing approaches realize this reasoning at the text level, leading to high latency and exposure bias from autoregressive decoding. Inspired by advances in latent-space reasoning for large language models [hao2024training], we shift this cognitive process to a unified latent space. Our reasoner performs that within only two forward passes, preserving the semantic richness of language-based reasoning while avoiding token-generation overhead. This interpretable design forms an efficient foundation for hierarchical parallel planning.
@@ 4 Experiments
DriveVLM [tian2024drivevlm] CoRL 2024 0.18 0.34 0.68 0.40
DriveVLM-Dual [tian2024drivevlm] CoRL 2024 0.15 0.29 0.48 0.31
OmniDrive [wang2024omnidrive] CVPR 2025 0.14 0.29 0.55 0.33 0.00 0.13 0.78 0.30
EMMA [hwang2024emma] TMLR 0.14 0.29 0.54 0.32
EMMA+ [hwang2024emma] TMLR 0.13 0.27 0.48 0.29
ImpromptuVLA [chi2025impromptu] NeurIPS 2025 0.13 0.27 0.53 0.30
SOLVE-VLM [Chen_2025_CVPR] CVPR 2025 0.13 0.25 0.47 0.28 0.00 0.16 0.43 0.20
@@
UniAD [hu2023planning] CVPR 2023 0.59 1.01 1.48 1.03 0.16 0.51 1.64 0.77
VAD-Base [jiang2023vad] ICCV 2023 0.69 1.22 1.83 1.25 0.06 0.68 2.52 1.09
BEV-Planner [li2024ego] CVPR 2024 0.30 0.52 0.83 0.55 0.10 0.37 1.30 0.59
UniAD [hu2023planning] CVPR 2023 0.20 0.42 0.75 0.46 0.02 0.25 0.84 0.37
VAD-Base [jiang2023vad] ICCV 2023 0.17 0.34 0.60 0.37 0.04 0.27 0.67 0.33
AD-MLP [admlp] arXiv 2023 0.15 0.32 0.59 0.35 0.00 0.27 0.85 0.37
BEV-Planner++ [li2024ego] CVPR 2024 0.16 0.32 0.57 0.35 0.00 0.29 0.73 0.34
SOLVE-E2E [Chen_2025_CVPR] CVPR 2025 0.14 0.28 0.50 0.31 0.04 0.17 0.68 0.30
@@
UniAD [hu2023planning] CVPR 2023 0.73 0.84 0.10 1.26 88.6 87.8 98.4 79.6
VAD [jiang2023vad] ICCV 2023 0.66 0.47 0.04 1.45 92.5 96.2 99.6 81.6
SparseDrive [sun2024sparsedrive] ICRA 2025 0.92 93.9
BridgeAD-S [zhang2025bridging] CVPR 2025 1.52 76.2
BridgeAD-B [zhang2025bridging] CVPR 2025 1.60 72.6
ImpromptuVLA [chi2025impromptu] NeurIPS 2025 2.06 2.55 1.86 1.78 65.1 54.8 72.8 67.6
BridgeAD-S [zhang2025bridging] CVPR 2025 2.98 46.1
BridgeAD-B [zhang2025bridging] CVPR 2025 3.06 44.3
@@
Table 2: Closed-loop simulation results on NeuroNCAP [ljungbergh2024neuroncap]. $\dagger$ indicates that it utilizes additional training data and is a text-based driving VLM model. $\ddagger$ refers to trajectory post-processing. Our proposed method achieves a substantial improvement in the closed-loop evaluation, demonstrating strong adaptability to safety-critical scenarios and highlighting the model’s efficiency and generalization capability. In this evaluation, we adopt only the top-1 driving strategy to better simulate realistic decision-making in closed-loop settings, ensuring fair comparison and faithfully reflecting the model’s safety and robustness in real-world driving situations. Best scores are in bold.
@@ 4.1 Experiment Settings
Datasets and evaluation metrics. We conduct experiments on the challenging nuScenes [caesar2020nuscenes] dataset, which contains 1,000 driving scenes of approximately 20 seconds each, providing six camera images, LiDAR data, semantic maps, and 3D bounding box annotations for keyframes. To enhance the planning-related supervision of our VLM, we additionally use the OmniDrive-nuScenes [wang2024omnidrive] extension, which augments nuScenes with QA-style annotations spanning perception, prediction, and planning. These annotations enable both offline reasoning supervision and online spatial reasoning.
We evaluate our model in both open-loop and closed-loop settings. Open-loop evaluation follows prior works [hu2023planning, jiang2023vad] and measures trajectory accuracy and safety using two standard metrics: the L2 displacement error (distance between predicted and ground-truth trajectories) and the collision rate. For closed-loop evaluation, we employ the NeuroNCAP simulator [ljungbergh2024neuroncap], a photorealistic framework built upon nuScenes that reconstructs diverse, safety-critical urban interactions for realistic policy assessment. Performance is measured by the NeuroNCAP Score, a five-star metric that rewards safe, low-impact behavior, and the collision rate, jointly reflecting both safety and operational efficiency. Together, these benchmarks provide a comprehensive evaluation of the model’s accuracy, robustness, and generalization in both static and dynamic driving scenarios.
Implementation details. Our implementation is based on the LLaVA v1.5 [liu2024visual] framework, which uses LLaMA-7B as the language model. Following prior work, we initialize the image encoder with EVA-02-L [fang2024eva] and adopt the SQ-Former [Chen_2025_CVPR] architecture for better visual reasoning. The detection decoder and lane decoder follow StreamPETR [wang2023exploring], with 900 object queries and 300 lane queries, respectively.
We use a multi-stage training strategy. In the first stage, our VLM is pretrained on QA pairs from OmniDrive-nuScenes, where LoRA layers [hu2021lora] are applied to prompt and adapt the VLM efficiently for perception–planning alignment. In the second stage, we integrate our action-based planner and perform joint fine-tuning. Within the VLM, we only update the LoRA parameters to retain pretrained knowledge while improving decision capability. Training is conducted with the AdamW optimizer [adamw] with Cosine Annealing [cosineanneal], using a weight decay of $1\times 10^{-4}$ and an initial learning rate of $1\times 10^{-4}$. Additional details and hyperparameters are provided in the supplementary material.
@@ 4.2 Comparison with State-of-the-art Methods
OmniDrive [wang2024omnidrive] 3727
SOLVE-VLM [Chen_2025_CVPR] 3719
@@
Table 3: Inference latency comparison on a single NVIDIA H20 GPU without flash-attention [dao2022flashattention]. We report end-to-end inference latency (ms per frame) under identical batch settings.
@@ 4.4 Ablation Studies
In this section, we systematically ablate the proposed components and key hyperparameters to assess their individual and combined contributions. We report results on nuScenes [caesar2020nuscenes] using two complementary metrics: the open-loop L2 average error and the closed-loop NeuroNCAP Score.

### v3

# arXiv:2512.22939v3[cs.CV] 27 Feb 2026

@@ 1 Introduction
Autonomous driving aims to predict safe and comfortable motion from rich multimodal observations. Early systems were organised as modular perception–prediction–planning stacks with dedicated 3D perception [37, 25, 40, 26, 55, 56, 36], forecasting [5, 15, 30, 16], and planning components [39, 13, 4, 10, 28]. Recent end-to-end approaches unify the stack and learn from pixels to waypoints or controls in a single pipeline [20, 24, 21, 7, 57, 48, 41]. In parallel, vision–language models (VLMs) [2, 49, 1, 47, 29] are increasingly integrated to inject cross-modal priors and world knowledge, either as agentic planning models that output controls or trajectories [52, 35, 38, 22, 50, 45] or as reasoning assistants that guide end-to-end modules [23, 42, 8].
Despite rapid progress, important gaps remain between accuracy in benchmarks and reliability in deployment. Modular systems provide interpretable components and strong geometric priors, but brittle interfaces can propagate errors and make global optimization difficult. End-to-end systems reduce manual interfaces and achieve high open-loop accuracy [20, 24, 21, 7, 57, 48], yet they often rely on sparse trajectory supervision, intertwine perception and control in ways that obscure causal structure, and struggle to generalize to out-of-distribution scenarios. Text-based VLM planners add powerful priors but introduce some practical issues: (1) Modality mismatch. Discrete text tokens do not align with the continuous geometry and dynamics of trajectories and can lead to format violations or physically inconsistent waypoints [35, 22]. (2) Chain-of-thought reasoning latency. The computation overhead of autoregressive decoding arises from its iterative token-by-token generation, where each new token depends on previously ones, causing the sequence to grow over time and significantly increasing inference delay.
@@ 2.1 End-to-end Autonomous Driving
Classical autonomous driving stacks decompose perception, prediction, and planning into separate modules, easing engineering but causing information loss and suboptimal coordination. End-to-end (E2E) systems instead learn a single differentiable mapping from sensor inputs to trajectories or control commands. Early demonstrations proved feasibility but suffered from limited transparency and data efficiency. Recent works enhance this paradigm with richer supervision and unified architectures: UniAD [21] integrates perception and planning within a Transformer, and VAD [24, 7] employs vectorized abstractions for efficiency and uncertainty modeling. Other directions explore task dependency [48], latent generative modeling [57, 28], or simplified ego-centric controllers [53, 27]. Despite progress, most E2E methods rely on sparse supervision from a single ground truth trajectory, which limits reasoning and diversity, highlighting the need for higher-level semantic and cognitive guidance.
@@ 2.2 VLMs for Autonomous Driving
Vision–Language Models (VLMs) bring broad world knowledge and remarkable reasoning ability to autonomous driving, inspiring two main integration paradigms. Single-system approaches treat planning as language modeling: visual inputs and prompts are translated into textual trajectories or control commands with chain-of-thought explanations [22, 34, 50, 52, 6]. Dual-system designs pair a VLM with an end-to-end planner, where the VLM proposes high-level commands or low-frequency trajectories to guide a lightweight action controller [42, 23, 8]. Other efforts align vision–language features in BEV [46], extend reasoning to 3D [45], or use VLMs as teachers to supervise E2E planning models [51]. While these methods improve interpretability and generalization, they still rely on explicit text-level reasoning or suffer from a feature gap between the VLM and the downstream action planner. This motivates our unified vision–language–action framework, which performs reasoning entirely in a shared latent space, bridging perception, cognition, and control within a single coherent architecture.
@@ 3.1 Framework Overview
Given multimodal representation $\mathbf{S}_{t}$ comprising multi–view images, LiDAR points, radar grids, a text prompt and ego state, or a subset thereof, the trajectory planning task reduces to mapping $\mathbf{S}_{t}$ to a $K$-step trajectory $\widehat{\mathbf{Y}}_{t}=[(x_{t+1},y_{t+1}),\ldots,(x_{t+K},y_{t+K})]$. Existing planning systems generally consist of two unified learnable components: Reasoner ($\mathcal{R}_{\theta}$), usually implemented as ResNet [18], PointNet [37], Q-Former [11], or VLM Transformer that extracts and fuses multimodal features or optionally refines them with chain-of-thought, and Planner ($\mathcal{P}_{\phi}$), usually instantiated as an deterministic MLP planner, stochastic diffusion sampler, or standard Transformer [43] decoder that injects action queries and regresses continuous way-points. Their interaction is succinctly captured by
@@ 3.2 Cognitive Latent Reasoning
Recent vision–language models show strong generalization in autonomous driving by performing a driver-like reasoning process: perceiving the scene, identifying critical entities, rethinking carefully and deciding final strategies. However, most existing approaches realize this reasoning at the text level, leading to high latency and exposure bias from autoregressive decoding. Inspired by advances in latent-space reasoning for large language models [17], we shift this cognitive process to a unified latent space. Our reasoner performs that within only two forward passes, preserving the semantic richness of language-based reasoning while avoiding token-generation overhead. This interpretable design forms an efficient foundation for hierarchical parallel planning.
@@ 4 Experiments
DriveVLM [42] CoRL 2024 0.18 0.34 0.68 0.40
DriveVLM-Dual [42] CoRL 2024 0.15 0.29 0.48 0.31
OmniDrive [45] CVPR 2025 0.14 0.29 0.55 0.33 0.00 0.13 0.78 0.30
EMMA [22] TMLR 0.14 0.29 0.54 0.32
EMMA+ [22] TMLR 0.13 0.27 0.48 0.29
ImpromptuVLA [9] NeurIPS 2025 0.13 0.27 0.53 0.30
SOLVE-VLM [8] CVPR 2025 0.13 0.25 0.47 0.28 0.00 0.16 0.43 0.20
@@
UniAD [21] CVPR 2023 0.59 1.01 1.48 1.03 0.16 0.51 1.64 0.77
VAD-Base [24] ICCV 2023 0.69 1.22 1.83 1.25 0.06 0.68 2.52 1.09
BEV-Planner [27] CVPR 2024 0.30 0.52 0.83 0.55 0.10 0.37 1.30 0.59
UniAD [21] CVPR 2023 0.20 0.42 0.75 0.46 0.02 0.25 0.84 0.37
VAD-Base [24] ICCV 2023 0.17 0.34 0.60 0.37 0.04 0.27 0.67 0.33
AD-MLP [53] arXiv 2023 0.15 0.32 0.59 0.35 0.00 0.27 0.85 0.37
BEV-Planner++ [27] CVPR 2024 0.16 0.32 0.57 0.35 0.00 0.29 0.73 0.34
SOLVE-E2E [8] CVPR 2025 0.14 0.28 0.50 0.31 0.04 0.17 0.68 0.30
@@
UniAD [21] CVPR 2023 0.73 0.84 0.10 1.26 88.6 87.8 98.4 79.6
VAD [24] ICCV 2023 0.66 0.47 0.04 1.45 92.5 96.2 99.6 81.6
SparseDrive [41] ICRA 2025 0.92 93.9
BridgeAD-S [54] CVPR 2025 1.52 76.2
BridgeAD-B [54] CVPR 2025 1.60 72.6
ImpromptuVLA [9] NeurIPS 2025 2.06 2.55 1.86 1.78 65.1 54.8 72.8 67.6
BridgeAD-S [54] CVPR 2025 2.98 46.1
BridgeAD-B [54] CVPR 2025 3.06 44.3
@@
Table 2: Closed-loop simulation results on NeuroNCAP [31]. $\dagger$ indicates that it utilizes additional training data and is a text-based driving VLM model. $\ddagger$ refers to trajectory post-processing. Our proposed method achieves a substantial improvement in the closed-loop evaluation, demonstrating strong adaptability to safety-critical scenarios and highlighting the model’s efficiency and generalization capability. In this evaluation, we adopt only the top-1 driving strategy to better simulate realistic decision-making in closed-loop settings, ensuring fair comparison and faithfully reflecting the model’s safety and robustness in real-world driving situations. Best scores are in bold.
@@ 4.1 Experiment Settings
Datasets and evaluation metrics. We conduct experiments on the challenging nuScenes [3] dataset, which contains 1,000 driving scenes of approximately 20 seconds each, providing six camera images, LiDAR data, semantic maps, and 3D bounding box annotations for keyframes. To enhance the planning-related supervision of our VLM, we additionally use the OmniDrive-nuScenes [45] extension, which augments nuScenes with QA-style annotations spanning perception, prediction, and planning. These annotations enable both offline reasoning supervision and online spatial reasoning.
We evaluate our model in both open-loop and closed-loop settings. Open-loop evaluation follows prior works [21, 24] and measures trajectory accuracy and safety using two standard metrics: the L2 displacement error (distance between predicted and ground-truth trajectories) and the collision rate. For closed-loop evaluation, we employ the NeuroNCAP simulator [31], a photorealistic framework built upon nuScenes that reconstructs diverse, safety-critical urban interactions for realistic policy assessment. Performance is measured by the NeuroNCAP Score, a five-star metric that rewards safe, low-impact behavior, and the collision rate, jointly reflecting both safety and operational efficiency. Together, these benchmarks provide a comprehensive evaluation of the model’s accuracy, robustness, and generalization in both static and dynamic driving scenarios.
Implementation details. Our implementation is based on the LLaVA v1.5 [29] framework, which uses LLaMA-7B as the language model. Following prior work, we initialize the image encoder with EVA-02-L [14] and adopt the SQ-Former [8] architecture for better visual reasoning. The detection decoder and lane decoder follow StreamPETR [44], with 900 object queries and 300 lane queries, respectively.
We use a multi-stage training strategy. In the first stage, our VLM is pretrained on QA pairs from OmniDrive-nuScenes, where LoRA layers [19] are applied to prompt and adapt the VLM efficiently for perception–planning alignment. In the second stage, we integrate our action-based planner and perform joint fine-tuning. Within the VLM, we only update the LoRA parameters to retain pretrained knowledge while improving decision capability. Training is conducted with the AdamW optimizer [32] with Cosine Annealing [33], using a weight decay of $1\times 10^{-4}$ and an initial learning rate of $1\times 10^{-4}$. Additional details and hyperparameters are provided in the supplementary material.
@@ 4.2 Comparison with State-of-the-art Methods
OmniDrive [45] 3727
SOLVE-VLM [8] 3719
@@
Table 3: Inference latency comparison on a single NVIDIA H20 GPU without flash-attention [12]. We report end-to-end inference latency (ms per frame) under identical batch settings.
@@ 4.4 Ablation Studies
In this section, we systematically ablate the proposed components and key hyperparameters to assess their individual and combined contributions. We report results on nuScenes [3] using two complementary metrics: the open-loop L2 average error and the closed-loop NeuroNCAP Score.
@@ 5 Conclusion
Supplementary Material
6 Further Ablation Study
Ablation on context tokens for planning. We further study whether the planner should attend to all visual tokens or only to the pruned subset selected by the ego-adaptive router. Concretely, we compare two variants: one that uses the full set of original vision tokens when predicting the final trajectory, and one that relies solely on the pruned critical tokens $\mathbf{Q}^{*}$ while keeping all other components unchanged. As summarized in Tab. 8, using pruned tokens leads to lower L2 errors compared to using all tokens. This indicates that the ego-adaptive router can accurately identify which visual cues are most relevant to the current driving scenario and future decisions, effectively filtering out redundant background information. At the same time, pruning substantially reduces the token length fed into the planner, thereby lowering computation and improving inference efficiency without sacrificing final prediction accuracy.
Visual tokens L2 (cm) $\downarrow$
1s 2s 3s Avg
Full tokens 14.8 28.9 52.4 32.0
Pruned tokens (Ours) 14.0 27.1 50.2 30.4
Table 8: Ablation on planner context token selection. Using the pruned critical tokens as context for our planner yields better accuracy while reducing sequence length and computational cost.
7 Further Implementation Details
Architecture. Our implementation is built on the LLaVA v1.5 framework [29], which uses LLaMA-7B as the language backbone and introduces LoRA adapters [19] for efficient adaptation. The image encoder follows EVA-02-L [14] instantiated as an EVAViT backbone with 24 transformer layers, hidden dimension 1024 and window size 16, initialized from publicly available weights. On top of the image backbone, we adopt the SQ-Former architecture [8] for multi-view visual reasoning and use perception task heads. We use 900 object queries and 300 lane queries by default, consistent with our main experiments. The overall system is implemented in a single camera-only configuration with 6 multi-view images per frame. And point-cloud-style BEV parameters (e.g., point cloud range and voxel size) are used only for defining the spatial grid for 3D perception heads.
Multi-stage training strategy. We employ a three-stage training pipeline to fully exploit the strong VLM priors while stabilizing downstream hierarchical planning:
  • Stage 1: VLM adaptation on OmniDrive QA. We first adapt the vision–language backbone on OmniDrive-nuScenes QA pairs [45]. Similar to OmniDrive, we use scene-level QA covering description, perception, prediction and planning questions, and optimize only the LoRA layers on top of the LLaMA-7B language model together with the Q-Former style visual interface. This stage encourages the VLM to encode driving-relevant semantics and reasoning skills in complex urban scenes without changing the base model weights, and aligns the visual encoder with driving-oriented questions.
  • Stage 2: Meta-action bank pretraining. In the second stage, we introduce the meta-action bank used by the Cognitive Latent Reasoner. We cluster nuScenes trajectories (using the provided $k$-means anchors in the config) to obtain a discrete set of meta-action prototypes, which are used to initialize the meta-action queries and the action bank. During this stage, we freeze the language backbone VLM (except LoRA), and primarily train the meta-action classification and trajectory prediction heads to associate latent meta-actions with typical driving patterns (e.g., straight, lane change, turn, braking). This provides a structured and stable prior for downstream hierarchical planning.
  • Stage 3: End-to-end fine-tuning on nuScenes. Finally, we fine-tune ColaVLA on nuScenes planning supervision. The language backbone remains frozen while LoRA adapters, the visual encoder (EVAViT + SQ-Former), detection/map heads, cognitive latent reasoning components (router and meta-action queries), and the hierarchical parallel planner are all jointly optimized. We train with multi-scale trajectory losses and the top-$K$ hypothesis strategy, which enables confidence-guided selection among multiple meta-action hypotheses.
Optimization and training schedule. Unless otherwise specified, all stages are trained on 16 NVIDIA H20 GPUs with a per-GPU batch size of 2 (effective batch size 32). We use the AdamW optimizer [32] with initial learning rate $1\times 10^{-4}$, weight decay $1\times 10^{-4}$, and $(\beta_{1},\beta_{2})=(0.9,0.999)$. The learning rate is scheduled by Cosine Annealing [33] with linear warmup for the first 500 iterations and a minimum learning-rate ratio of $10^{-3}$. We train for 10 epochs on nuScenes dataset in the final stage, using mixed-precision (FP16) with dynamic loss scaling and gradient clipping at a global norm of 10.
Data processing. We follow the standard nuScenes setting [3] with the official train/val split. Our data loader uses CustomNuScenesDataset with temporal sequences, camera-only inputs, and the same object classes as the main experiments. Images are first loaded at $900\times 1600$ resolution and then augmented with random resize–crop–rotation as in the provided augmentation configure. For training, we resize each view to $320\times 640$ and normalize them using ImageNet statistics. We adopt the same pipeline for validation but disable random augmentation. Planning supervision is derived from nuScenes ego trajectories, while QA supervision and lane-object relations are loaded from the OmniDrive and preprocessed lane-object files.
Summary of key hyperparameters. For clarity, Tab. 9 summarizes the main architectural and training hyperparameters used in our implementation.
Category Setting
Backbone VLM LLaVA v1.5 (LLaMA-7B) with LoRA adapters
Image encoder EVA-02-L (EVAViT, 24 layers, 1024-d, window size 16)
Visual reasoning SQ-Former with temporal PETR-style transformers
Object queries 900 queries (3D detection head)
Lane queries 300 queries (map / lane head)
Input modality 6-camera views, camera-only, no LiDAR or radar
Image resolution $900\times 1600\rightarrow 320\times 640$ after augmentation
Point cloud range [51.2,51.2,5.0, 51.2,51.2,3.0][-51.2,-51.2,-5.0,\;51.2,51.2,3.0]
Voxel size $[0.2,0.2,8.0]$ (for BEV grid definition)
Planner stages $S=6$ hierarchical scales (as in stage_index)
Multi-scale loss Weights $[0.5,0.7,1.0,1.2,1.5,1.8]$
Top-$K$ hypotheses $K=3$ meta-action modes (topk_mode_predict=3)
Batch size 2 samples / GPU ($\times 16$ GPUs)
Epochs (stage 3) 10 epochs on nuScenes
Optimizer AdamW, lr $1\times 10^{-4}$, wd $1\times 10^{-4}$, $(0.9,0.999)$
LR schedule Cosine Annealing + 500-iter linear warmup, min lr ratio $10^{-3}$
Precision FP16 with dynamic loss scaling
Grad clipping Global norm 10
Efficiency tricks Gradient checkpointing, FlashAttention, mixed precision
Table 9: Main implementation and training hyperparameters for ColaVLA.
Figure 5: Qualitative closed-loop comparisons in the NeuroNCAP simulator across three representative scenario types. For each case, we visualize the predicted trajectories of ColaVLA and competing planners. ColaVLA consistently guides the ego vehicle away from potential collisions, producing safer and more stable motions.
8 Additional Closed-Loop Visualizations
We further provide qualitative results from the NeuroNCAP simulator to complement the nuScenes open-loop visualizations in the main paper. As illustrated in Fig. 5, our model produces stable and collision-averse behaviors across diverse safety-critical scenarios, including static obstacles, frontal interactions, and side conflicts. Compared with previous text-based VLM planners, ColaVLA exhibits more responsive closed-loop reactions and fewer failure cases, while simultaneously maintaining higher inference efficiency. These examples corroborate our quantitative findings that ColaVLA achieves stronger NeuroNCAP scores and lower latency, delivering both safer and faster closed-loop planning.
9 Details about Meta-Action Bank
To support cognitive latent reasoning, we pre-compute a discrete meta-action label for each nuScenes sample and use it to index the action bank. Each meta-action is defined by combining a path pattern with a speed pattern derived jointly from the future ego trajectory and CAN bus signals. In our final setup, this yields eight meta-action classes: four straight-driving modes (stopping, constant-speed cruising, accelerating, and decelerating along a roughly straight path) and four turning or lane-change modes (large left turns or U-turns, large right turns or U-turns, small left turns or lane changes, and small right turns or lane changes). Our implementation broadly follows the trajectory categorization procedure used in Senna [23].
Geometry and CAN bus fusion. For each sample, we first obtain a future ego trajectory in the ego frame as a sequence of $T$ waypoints at a fixed time interval. If the planning trajectory is missing or fully padded, we reconstruct it from nuScenes CAN bus pose messages by transforming global positions into the initial ego frame and resampling to $T$ steps; if this is still not possible, we fall back to a zero trajectory. In parallel, we read CAN bus logs around the planning horizon (pose, yaw, longitudinal velocity and acceleration), and compute fused heading and speed statistics such as cumulative yaw angle, maximum yaw rate, final speed, speed change, and mean acceleration. When available, CAN bus statistics are preferred; otherwise, we rely on geometric estimates from the trajectory itself.
Path classification. We first assign each sample to one of five path types: straight motion, large left turn (including left U-turns), large right turn (including right U-turns), small left turn or lane change, and small right turn or lane change. The assignment is determined using fused heading and lateral-motion statistics. Large turns are identified by large cumulative yaw with a clear left or right sign; small turns correspond to moderate yaw changes; lane-change-like motions are characterized by small net yaw but significant lateral displacement with consistent lateral direction and bounded yaw rate; straight motion is defined by small yaw and small lateral offset. Ambiguous residual cases are assigned to left or right small turns according to the sign of the lateral displacement. This step yields a path label for every sample.
Speed classification and heavy paths. Speed patterns are only refined for path types that appear frequently in the training set. We first count occurrences of each path type and mark heavy paths whose counts exceed a global threshold. In practice, only straight paths are heavy, while turning and lane-change paths are relatively sparse.
For heavy paths, we further classify the speed profile into four modes: stopping, approximately constant speed, accelerating, and decelerating. The decision uses fused speed statistics: very low final speed indicates stopping; clearly increasing speed or positive mean acceleration indicates acceleration; clearly decreasing speed or negative mean acceleration indicates deceleration; otherwise the sample is treated as constant-speed. For non-heavy path types, we do not split by speed and simply mark the speed pattern as “any”, in order to avoid severe class imbalance.
Final meta-action labels. The final meta-action class for each sample is defined as the combination of its path type and speed type. After applying the heavy-path and speed rules above, only eight combinations are active in training: four straight-driving modes (stop, constant-speed, accelerate, decelerate along a straight path) and four turning or lane-change modes (large left, large right, small left, small right). Each combination is assigned a stable integer id, which we precompute offline and store alongside each nuScenes sample. During training, the cognitive latent reasoner predicts a distribution over these eight meta-actions, and the top-scoring modes are used to retrieve corresponding action priors from the action bank, providing structured, trajectory-aligned guidance for the hierarchical parallel planner.